iosangel:

find the resultant of vector A-B where A =(5,-5) and B=(3,-8)

antman1:

Do you know the answers

antman1:

like what are you trying to figure out to find the answers

iosangel:

iosangel:

Options?

antman1:

well what are the answer options do you think?

iosangel:

I'm not sure

antman1:

well did u guess

iosangel:

I'm not sure where to start

antman1:

well i think u need to show your work

Vocaloid:

For A-B subtract the coordinates of A and B Since A is (5,-5) and B=(3,-8), A - B = (5 - 3, -5 - [-8])